# Thermal Stoner-Wohlfarth Model for Magnetodynamics of Single Domain Nanoparticles: Implementation and Validation

### v1

@@ III Conclusions
Driven by a long-standing need to introduce the internal magnetization dynamics into coarse-grained computer simulations of magnetic soft matter, we proposed an approach that is able to encompass both Neél and Brownian relaxations in a computationally affordable manner. In this paper, we introduce the Stoner-Wohlfarth thermal model (tSW) and validate it against an ensemble of immobilized, non-interacting, randomly oriented uniaxial particles (solid superparamagnet) and a classical dilute (non-interacting) ferrofluid for various combinations of anisotropy strength and Zeeman energy at a fixed temperature. In order to find the range of validity of our approach, we compared its predictions to the analytical results. The availability of the latter, in fact, defined the choice of the systems to model for this work.
It was crucial to verify whether the newly formulated tSW model is able to reproduce both the equilibrium and dynamic properties of magnetic soft matter systems. Our findings indicate that the tSW model is applicable and efficient when the anisotropy of a particle exceeds thermal fluctuations by a factor of four to five and above.
Our work opens possibilities for future investigations of magnetic soft matter systems in which the interplay between internal magnetization degrees of freedom and Brownian rotation is often overlooked.
@@ IV.1 Simulation Method
We perform MD simulations of a solid superparamagnet and a non-interacting ferrofluid for different anisotropy and Langevin parameter combinations $(\sigma,\xi)$, at a fixed temperature, using the ESPResSo simulation package [68]. Magnetic nanoparticles are assumed to be mono-disperse, spherical particles with a point dipole, and a characteristic diameter $\Omega$. The solvent was implicitly represented via the Langevin thermostat[69]. The equations of motion are integrated over time $t$ numerically:
@@
The performance and reliability of the implementation of tSW is dependent on a robust and highly efficient strategy to find the extrema in Eq.5. For this purpose, we make use of Nlopt, a free/open-source library for nonlinear optimization [72]. Nlopt is a versatile library written in C, with a common interface for a multitude of different optimization strategies. We have specifically settled on using the Method of Moving Asymptotes (MMA) algorithm [73] to find extrema in Eq.5. MMA is guaranteed to converge to some local minimum and has, throughout our testing, proven to be well suited for optimisation of Eq.5, converging to the correct energy minima very quickly.
@@ IV.2 Reduced units and mapping to physical parameters
In this subsection, we give an overview of the SI and reduced units used in our simulations. The units are presented in the order we find most instructive. We start with a common reference material choice for the magnetic nanoparticles, namely, magnetite. We assumed that magnetite has uniaxial anisotropy with a magnetic anisotropy constant of $K=10\;kJ/m^{3}$, saturation magnetisation $M_{s}=480\cdot 10^{3}\;kg/(ms^{2})$ and a core density $\rho_{m}=5170kg/m^{3}$. The uniaxial anisotropy assumption is commonly used, but is not strictly correct, as pointed out in Witt et al. [74]. In all cases, we assumed that the magnetite core is coated with a 2 nm thick oleic acid coating (density $895kg/m^{3}$). Given a value of $\sigma$ we want to explore, considering the material values given above, the unit length is chosen to be the diameter of the nanoparticle and a unit mass is chosen to be its mass. We choose the particle volume fraction in the simulation box to be $\phi=0.001$. The side length of the simulation box was set to $\approx 64\Omega$, derived based on the chosen $\phi$ and the particle number. In all cases, we choose the unit energy to be room temperature $T=298.15K$. In other words, we set the reduced temperature of the Langevin thermostat to $k_{B}T=1$. These three choices also uniquely determine the unit time. From there, we calculate $\tau_{D}$, where the gyromagnetic ratio is $\mu_{0}\cdot(1.76\cdot 10^{11}\;s^{-1}T^{-1})$, the damping parameter is chosen to be $\alpha=0.08$ and $\mu_{0}$ is the vacuum permittivity. We choose $\tau_{D}/\tau_{B}=0.01$. Based on this choice, we can derive the density of the implicit fluid and the respective rotational and translational friction coefficients.
@@ Acknowledgements
This research has been supported by the Project SAM P 33748. Computer simulations were performed at the Vienna Scientific Cluster (VSC-5). SSK acknowledges the support of Project DEMMON PAT 4120124, and DN MAESTRI (Marie Skłodowska-Curie Actions-Doctoral Networks grant agreement No 101119614).
@@ V Appendix A
Here, we derive the equilibrium magnetisation for an ensemble of immobilized, non-interacting, randomly oriented uniaxial particles. First, let us consider immobilized particles with a predefined orientation of the anisotropy axes ($\vec{n}=const$).The equilibrium projection of magnetization on the field direction is

### v2

@@ III Conclusions
Driven by a long-standing need to incorporate internal magnetization dynamics into coarse-grained computer simulations of magnetic soft matter, we proposed an approach capable of encompassing both Néel and Brownian relaxations in a computationally affordable manner. In this paper, we introduced the Stoner-Wohlfarth thermal model (tSW) and validated it against an ensemble of immobilized, non-interacting, randomly oriented uniaxial particles (solid superparamagnet) and a classical dilute (non-interacting) ferrofluid for various combinations of anisotropy strength and Zeeman energy at a fixed temperature. The choice of systems was dictated by the availability of analytical results, which were used to establish the range of validity of our approach. It was crucial to verify whether and when the tSW model could reproduce both the equilibrium and dynamic properties of magnetic soft matter systems. Our findings indicate that the tSW model is applicable and efficient when the anisotropy of a particle exceeds thermal fluctuations by a factor of five or more. We also applied the tSW model to interacting systems, uncovering and rationalizing a pronounced magnetization decrease in solid superparamagnets due to the dipole-interaction-induced formation of magnetic vortices/holes. The linear dependence of the proportion of these holes on $\xi$ fits most of our data, though its physical meaning and general functional form for arbitrary $\xi$ require further investigation. This work opens possibilities for future studies of magnetic soft matter systems, where the interplay between Néel and Brownian relaxation is often overlooked. We plan to address field-dependent relaxation dynamics within the tSW model in a future study.
@@ IV.1 Simulation Method
We perform MD simulations of a solid superparamagnet and a non-interacting ferrofluid for different anisotropy and Langevin parameter combinations $(\sigma,\xi)$, at a fixed temperature, using the ESPResSo simulation package [81]. Magnetic nanoparticles are assumed to be mono-disperse, spherical particles with a point dipole, and a characteristic diameter $d$. The solvent was implicitly represented via the Langevin thermostat [82]. The equations of motion are integrated over time $t$ numerically:
@@
The long-range dipole-dipole pair interactions are simulated using direct summation with two replicas. The performance and reliability of the implementation of tSW is dependent on a robust and highly efficient strategy to find the extrema in Eq.5. For this purpose, we make use of Nlopt, a free/open-source library for nonlinear optimization [85]. Nlopt is a versatile library written in C, with a common interface for a multitude of different optimization strategies. We have specifically settled on using the Method of Moving Asymptotes (MMA) algorithm [86] to find extrema in Eq.5. MMA is guaranteed to converge to some local minimum and has, throughout our testing, proven to be well suited for optimisation of Eq.5, converging to the correct energy minima very quickly.
@@ IV.2 Reduced units and mapping to physical parameters
In this subsection, we give an overview of the SI and reduced units used in our simulations. The units are presented in the order we find most instructive. We start with a common reference material choice for the magnetic nanoparticles, namely, magnetite. We assumed that magnetite has uniaxial anisotropy with a magnetic anisotropy constant of $K=10\;kJ/m^{3}$, saturation magnetization $M_{s}=480\cdot 10^{3}\;A/m$ and a core density $\rho_{m}=5170\;kg/m^{3}$. The uniaxial anisotropy assumption is commonly used, but is not strictly correct, as pointed out in Witt et al. [87]. In all cases, we assumed that the magnetite core is coated with a 2 nm thick oleic acid coating (density $895kg/m^{3}$). Given a value of $\sigma$ we want to explore, considering the material values given above, the unit length is chosen to be the diameter of the nanoparticle, $d$, and a unit mass is chosen to be its mass. We choose the particle volume fraction in the simulation box to be $\phi=0.001$. The side length of the simulation box was set to $\approx 64d$, derived based on the chosen $\phi$ and the particle number, $N$. In all cases, we choose the unit energy to be room temperature $T=298.15K$. In other words, we set the reduced temperature of the Langevin thermostat to $k_{B}T=1$. These three choices also uniquely determine the unit time. From there, we calculate $\tau_{D}$, where the gyromagnetic ratio is $\mu_{0}\cdot(1.76\cdot 10^{11}\;s^{-1}T^{-1})$, the damping parameter is chosen to be $\alpha=0.08$ and $\mu_{0}$ is the vacuum permittivity. We choose $\tau_{D}/\tau_{B}=0.01$. Based on this choice, we can derive the density of the implicit fluid and the respective rotational and translational friction coefficients.
@@ Acknowledgements
This research has been supported by the Project SAM P 33748. Computer simulations were performed at the Vienna Scientific Cluster (VSC-5). SSK acknowledges the support of Project DEMMON PAT 4120124, and DN MAESTRI (Marie Skłodowska-Curie Actions-Doctoral Networks grant agreement No 101119614). The authors thank Prof. Alexey O. Ivanov for guidance and feedback during the development of this work.
@@ V Appendix A
Here, we derive the equilibrium magnetization for an ensemble of immobilized, non-interacting, randomly oriented uniaxial particles. First, let us consider immobilized particles with a predefined orientation of the anisotropy axes ($\vec{n}=const$).The equilibrium projection of magnetization on the field direction is